OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

@fateal

OpenStudy (fateal):

sorry idk how to do it :/

OpenStudy (anonymous):

lol same

OpenStudy (anonymous):

@LichKing

OpenStudy (fateal):

i just forgot how to plug in

OpenStudy (kkutie7):

use the points to find the slope first using: \[\frac{y2-y1}{x2-x1}\]

OpenStudy (kkutie7):

point slope form is this: y+y1=m(x-x1)

OpenStudy (kkutie7):

x1=-3, x2=2, y1=2, y2=1

OpenStudy (kkutie7):

in the slope form equation y and x aren't values so no need to plug in there. x1 and y1 are values. m is the slope using the equation above. if you need more help I'd be happy to walk you through.

OpenStudy (anonymous):

so it is c right?